A communication bus has a Bit Error Rate (BER) of 10%. If you send a 3-bit message what is the probability that the message is received with at least one error?
Answer: B
Explanation: It is easier to calculate the probability of no errors and subtract from 1. P(no error) = (0.9)^3 = 0.729. Therefore, P(at least one error) = 1 - 0.729 = 0.271 or 27.1%.